### v1

# Quantum correction to the orbital Hall effect

Abstract
The orbital Hall effect (OHE) typically is evaluated using an intuitive approach based on semiclassical considerations. Here we provide a full quantum mechanical evaluation of the orbital current, including all matrix elements of the position operator. We recover previous results and find an additional quantum correction due to the non-commutativity of the position and velocity operators. We evaluate the full OHE for the topological antiferromagnet CuMnAs showing that the quantum correction dominates the response and changes its sign. We discuss implications for experiment.
Introduction. The non-equilibrium properties of Bloch electrons’ orbital angular momentum (OAM) Yafet (1963); Vanderbilt (2018) have come under intense scrutiny with the advent of orbitronics Bernevig et al. (2005); Das (2023); Rhonald Burgos Atencia and Culcer (2024), whose focus is generating non-equilibrium OAM densities and currents Choi et al. (2023); Liu and Culcer (2024); Xiao et al. (2021); Pezo et al. (2022); Salemi and Oppeneer (2022); Cysne et al. (2021); Luis M. Canonico (2024); Go et al. (2018); Jo et al. (2018); Sala and Gambardella (2022a); Wang et al. (2023a); Lyalin et al. (2023); Dutta and Tulapurkar (2022); Sala et al. (2023); Zhang et al. (2022); Ding et al. (2022); Voss et al. (2024). A significant technological motivation behind recent efforts is the electrical manipulation of magnetic degrees of freedom Kim et al. (2021); Chen et al. (2018); Go and Lee (2020); Lee et al. (2021a); Zheng et al. (2020); Lee et al. (2021b, c); Ding et al. (2020); Sala and Gambardella (2022b); Hayashi et al. (2023); Li et al. (2023), with an emphasis on weakly spin-orbit coupled materials Salvador-Sánchez et al. (2024); Seifert et al. (2023); Ünzelmann et al. (2020); Salemi et al. (2019); Tang and Bauer (2024); Wang et al. (2023b); Busch et al. (2024). From a fundamental science perspective, the OHE has been proposed as one of the mechanisms behind the observed valley Hall effect Bhowal and Vignale (2021); Wu et al. (2013); Song et al. (2015); Wu et al. (2013); Bhowal and Satpathy (2020); Xue et al. (2020), and attempts to disentangle these two effects are now underway Cysne et al. (2021, 2022); Kazantsev et al. (2024).
Determining the OAM and OHE involves working with the position operator, which is challenging in extended systems Fonseca et al. (2023); Ado et al. (2024); Yoda et al. (2018, 2015); Park et al. (2012, 2013, 2011). The main obstacle is the band-diagonal elements of the position operator. This obstacle can be circumvented in equilibrium, since the equilibrium OAM involves only inter-band matrix elements of the position operator. At present, the approach universally adopted for non-equilibrium systems is to start from the equilibrium OAM matrix elements and combine those with a non-equilibrium distribution found using standard methods such as the Boltzmann equation or the Kubo formula Rhonald Burgos Atencia and Culcer (2024); Liu and Culcer (2024); Cysne et al. (2022); Bhowal and Vignale (2021); Cysne et al. (2021); Pezo et al. (2022, 2023); Hanke et al. (2016); Go et al. (2020, 2018); Jo et al. (2018); Baek and Lee (2021); Kim et al. (2021); Tang and Bauer (2024); Inoue et al. (2004, 2006); Chen et al. (2024); Costa et al. (2023); Barbosa et al. (2023). This includes earlier work by some of us Liu and Culcer (2024). We argue that this approach, which has never been justified on fundamental grounds, is incomplete and misses important terms, a fact that can be seen already at the operator level. In the modern theory of orbital magnetization Vanderbilt (2018) the OAM is expressed as the expectation value of the symmetrized vector product of the position and velocity operators. To be consistent with the modern theory, the orbital current needs to be evaluated as the full quantum mechanical expectation value of the product of the OAM and velocity operators, including the band-diagonal terms in the position operator, and this needs to be found explicitly for a non-equilibrium system.
Figure 1: The intrinsic OHE for CuMnAs [Eq. 6] with OAM polarization along $\hat{\bm{z}}$ and electric field along $\hat{\bm{x}}$. The orbital current $j^{\vphantom{\dagger}}_{L}$ is given by the conventional contribution $j_{\rm conv}$ plus the quantum correction $\Delta j$. The parameters are $t=0.08\,{\rm eV}$ and $\tilde{t}=1\,{\rm eV}$, $\alpha_{\rm R}=0.8,\alpha_{\rm D}=0$,and ${\bm{h}}_{\rm AFM}=(0.85,0,0)\,{\rm eV}$.
In this paper we present a full quantum mechanical evaluation of the non-equilibrium expectation value of the orbital current operator, including all matrix elements of the position operator. Our principal finding is the full expression for the orbital current in response to an electric field, which is given by $j^{\vphantom{\dagger}}_{L}=j_{\rm conv}+\Delta j$ and is shown in Fig. 1 for CuMnAs. Here $j_{\rm conv}$ is the conventional orbital current, which has been evaluated to date, and is based on the equilibrium OAM and the non-equilibrium distribution. The quantum correction $\Delta j$ consists of contributions from: (i) the electron group velocity leading to band off-diagonal matrix elements of the OAM; (ii) inter-band coherence terms induced by the band-diagonal part of the position operator; and (iii) the non-commutativity of the position and velocity operators. We develop a quantum kinetic equation for the density matrix and evaluate the full OHE for the tight-binding band structure of the topological antiferromagnet CuMnAs, where $\Delta j$ provides the dominant contribution to the OHE, such that the resulting OHE has the opposite sign to the conventional formula. Our result is general and applies to all solids, having important implications for the interpretation of experiments. We focus on the intrinsic case and in closing discuss generalizations to disordered systems under certain approximations. We stress that our concern is with the evaluation of the orbital current, not its definition, which we take to be the usual product of the OAM and velocity operators.
OAM and orbital current operators. We begin by considering a generalized dipole operator $\hat{\bm{d}}^{A}={\textstyle{1\over 2}}\{\hat{A},\hat{\bm{r}}\}$, where $\hat{A}$ represents a general operator which may be a scalar such as charge, a vector component of spin, velocity, etc. The expectation value of such an operator is Tr$(\hat{\bm{d}}^{A}\hat{\rho})$ with $\hat{\rho}$ the density matrix. We define the OAM operator as the symmetrized combination $\hat{\bm{L}}={\textstyle{1\over 2}}(\hat{\bm{r}}\times\hat{\bm{v}}+\hat{\bm{v}}\times\hat{\bm{r}})$, with $\hat{\bm{v}}$ the velocity operator, and we work in the Hilbert space spanned by Bloch wave-functions ${\big{|}\hskip 1.00006pt\Psi_{m{\bm{k}}}\hskip 1.00006pt\big{\rangle}}=e^{i{\bm{k}}\cdot{\bm{r}}}{\big{|}\hskip 1.00006ptu_{m{\bm{k}}}\hskip 1.00006pt\big{\rangle}}$. Determining the OAM expectation value is tantamount to substituting the appropriate velocity component for $\hat{A}$ above, whereupon the correct expression for the equilibrium OAM in the modern theory immediately emerges.
The orbital current operator is ${\hat{\jmath}}^{\alpha}_{\delta}={\textstyle{1\over 2}}\big{\{}{\hat{L}}_{\alpha},{\hat{v}}_{\delta}\big{\}}$ and its expectation value is $\mathop{\textsf{Tr}}\,({\hat{\jmath}}^{\alpha}_{\delta}\,{\hat{\rho}})$. Using the cyclic property of the trace we decompose the orbital current operator into a dipole-like part and a commutator part as follows:
$j^{\alpha}_{\delta}={\textstyle{1\over 8}}\epsilon_{\alpha\beta\gamma}\mathop{\textsf{Tr}}\Big{(}\big{\{}\{{\hat{r}}_{\beta},\hat{\rho}\},\{{\hat{v}}_{\gamma},{\hat{v}}_{\delta}\}\big{\}}+\big{\{}{\hat{\rho}},\big{[}{\hat{v}}_{\gamma},[{\hat{r}}_{\beta},{\hat{v}}_{\delta}]\big{]}\big{\}}\Big{)}\ .$ (1)
This decomposition helps fix the conventional orbital current and shed light onto the additional terms. The first term contains the conventional orbital current $j_{\rm conv}$. This can be obtained immediately by retaining only the inter-band terms in ${\hat{r}}_{\beta}$, ${\hat{v}}_{\gamma}$ and ${\hat{v}}_{\delta}$. The quantum correction stems from: (i) the band-diagonal terms in the velocity operators; (ii) the band-diagonal terms in the position operator, and (iii) the commutator $[\hat{r}_{\beta},\hat{v}_{\delta}]$ in the second term, which reflects the non-commutativity of the position and velocity operators. In the crystal momentum representation the commutator matrix elements are
$i\big{[}{\hat{r}}_{\beta},{\hat{v}}_{\delta}\big{]}^{nm}_{\bm{k}}=\bigg{(}{{\partial}v_{\delta}\over{\partial}k_{\beta}}\bigg{)}^{\!\!nm}_{\!\!{\bm{k}}}-i\,\big{[}{\mathcal{R}}_{\beta},v_{\delta}\big{]}^{nm}_{\bm{k}}\equiv\bigg{(}{Dv_{\delta}\over Dk_{\beta}}\bigg{)}^{\!\!nm}_{\!\!{\bm{k}}}\ ,$ (2)
where ${\mathcal{R}}_{\beta}^{lm}=i\,{\langle\hskip 1.00006ptu_{l}\hskip 1.00006pt|\hskip 1.00006pt{{\partial}u_{m}\over{\partial}k_{\beta}}\hskip 1.00006pt\rangle}$ is the Berry connection, which reflects the wave vector dependence of the basis functions. The velocity operator in turn has matrix elements
$\hbar\,[{\hat{v}}_{\delta}]^{mn}_{\bm{k}}={{\partial}{\varepsilon}_{m}\over{\partial}k_{\delta}}\,\delta_{mn}+i\,{\mathcal{R}}^{mn}_{\delta}({\varepsilon}_{m}-{\varepsilon}_{n})\ ,$ (3)
where ${\varepsilon}_{m}={\varepsilon}_{m{\bm{k}}}$ is the band energy at wavevector ${\bm{k}}$.
To evaluate the full OHE we require the non-equilibrium correction to the density matrix in an electric field, for which we turn to linear response theory following the approach of Refs. Culcer et al. (2017); Atencia et al. (2022). The single-particle density operator obeys the quantum Liouville equation, ${\partial}{\hat{\rho}}/{\partial}t+(i/\hbar)[{\hat{H}},{\hat{\rho}}]=0$, where ${\hat{H}}={\hat{H}}_{0}+{\hat{H}}_{E}$. Here ${\hat{H}}_{0}$ is the band Hamiltonian and ${\hat{H}}_{E}=e{\bm{E}}\cdot\hat{\bm{r}}$ is potential due to the external electrical field. At this stage we focus on intrinsic effects and do not consider disorder scattering, which will be discussed in closing. In the crystal momentum representation the equilibrium density matrix has the diagonal form $\rho_{0{\bm{k}}}^{mn}=f_{m}\,\delta_{mn}$, where $f_{m}\equiv f({\varepsilon}_{m{\bm{k}}})$ is the Fermi-Dirac distribution for band $m$. In an electric field the density matrix can be written as $\hat{\rho}=\rho_{0}+\rho_{E}$, and, in linear response, it has been shown that Culcer et al. (2017)
[…]
Figure 2: The quantum correction $\Delta j$ is decomposed as $\Delta j=\Delta{j_{1}}+\Delta{j_{2}}+\Delta{j_{3}}$. $\Delta j_{1}$ is from the group velocity, $\Delta j_{2}$ is from the band-diagonal part of position operator, and $\Delta j_{3}$ is from the non-commutativity of the position and velocity operators. We have used the same parameters as in Fig. 1.
Once $\rho^{mn}_{E{\bm{k}}}$ is found the quantum correction can be written as $\Delta j=\Delta j_{1}+\Delta j_{2}+\Delta j_{3}$, where
$\Delta j^{\alpha}_{\delta,1}$ $={\textstyle{1\over 2}}\,\epsilon_{\alpha\beta\gamma}\sum_{m,{\bm{k}}}\big{\{}{\mathcal{R}}^{\vphantom{\dagger}}_{\beta},\rho^{\vphantom{\dagger}}_{E{\bm{k}}}\big{\}}^{mm}\,v^{mm}_{\delta}v^{mm}_{l}$ (5)
$\Delta j^{\alpha}_{\delta,2}$ $={\textstyle{i\over 4}}\,\epsilon^{\vphantom{*}}_{\alpha\beta\gamma}\!\sum_{m,n,{\bm{k}}}{2eE_{\mu}\left[{D\Xi^{0}_{\beta}\over Dk_{\mu}}\right]^{mn}\!\!\!+\big{\{}\hbar v^{\vphantom{\dagger}}_{\beta},\rho^{\vphantom{\dagger}}_{E{\bm{k}}}\big{\}}^{mn}\over{\varepsilon}_{n}-{\varepsilon}_{m}}\{v_{\gamma},v_{\delta}\}^{nm}$
$\Delta j^{\alpha}_{\delta,3}$ $={\textstyle{i\over 4}}\,\epsilon_{\alpha\beta\gamma}\sum_{m,n,{\bm{k}}}\Big{[}v_{\gamma},{Dv_{\delta}\over Dk_{\beta}}\Big{]}^{mn}_{\bm{k}}\rho^{nm}_{E{\bm{k}}}\quad,$
where we have abbreviated $\big{[}\Xi^{0}_{\beta}\big{]}^{mn}={\textstyle{1\over 2}}{\mathcal{R}}^{mn}_{\beta}(f_{m}+f_{n})$, and $m\neq n$ is understood in all the summations. The result is gauge invariant, as shown in the Supplement. The OAM polarization is taken to be along the $\alpha$-direction while the transport direction is denoted by $\delta$. The three terms entering the quantum correction reflect inter-band coherence induced by the electric field. $\Delta j_{1}$ has the same structure as $j_{\rm conv}$, except that the electron group velocity appears instead of the inter-band velocity. $\Delta j_{2}$ is the result of restoring the band-diagonal matrix elements of the position operator on the left side of Eq. 1. Finally, $\Delta j_{3}$ reflects the non-commutativity of the position and velocity operators, which will be discussed below.
Tetragonal CuMnAs. We now exemplify our evaluation by studying the OHE in a tight-binding model of CuMnAs. CuMnAs can exist in both orthorhombic and tetragonal crystal structures. Both have been identified experimentally Linn et al. (2023); Wadley et al. (2015), and both can host the antiferromagnetic Dirac semimetal phase. Here we take the stable tetragonal CuMnAs phase as an example, which has opposite spins lying on a bipartite lattice. Such an arrangement breaks the $\mathcal{P}$ as well as the $\mathcal{T}$ symmetry locally. However, the combined $\mathcal{PT}$ symmetry is preserved by the exchange of the sublattices with the flip of oppositely aligned spins Das et al. (2023); Watanabe and Yanase (2021a), enforcing the bands at each momentum is doubly degenerate Li et al. (2020). The tight-binding Hamiltonian for CuMnAs without external magnetization is written as Wang et al. (2021)
$H_{0{\bm{k}}}=\begin{pmatrix}{\varepsilon}_{0}({\bm{k}})+{\bm{h}}({\bm{k}})\cdot{\bm{\sigma}}&V_{\rm AB}({\bm{k}})\\ V_{\rm AB}({\bm{k}})&{\varepsilon}_{0}({\bm{k}})-{\bm{h}}({\bm{k}})\cdot{\bm{\sigma}}\end{pmatrix}\quad,$ (6)
Here ${\varepsilon}_{0}({\bm{k}})=-t(\cos k_{x}+\cos k_{y})$ and $V_{\text{AB}}=-2\tilde{t}\cos(k_{x}/2)\cos(k_{y}/2)$, where $t$ and $\tilde{t}$ denote hopping between orbitals of the same and different sublattices, respectively. The sublattice-dependent spin-orbit coupling and the magnetization field are included in
[…]
with $\alpha^{\vphantom{\dagger}}_{\rm R}$ and $\alpha^{\vphantom{\dagger}}_{\rm D}$ the Rashba and Dresselhaus type spin-orbit coupling coefficients Watanabe and Yanase (2021b). In the band dispersion gapless points appear along the high symmetry line $(k_{x}=\pi)$, and we consider the local topology around the high-symmetry point. We expand the Hamiltonian in the vicinity of $(k_{x},k_{y})=(1,0.5)\pi$ up to $\delta k_{x,y}=\pm 0.1\pi$. Then we numerically integrate the full OHE for the approximated model Hamiltonian around $(1,0.5)\pi$ up to ±$\pm 0.1\pi$ by ${\bm{k}}$-discretization. The intrinsic anomalous Hall effect is proportional to the integration over the Fermi sea of the Berry curvature of each occupied band or equivalently to the integral of the Berry phases over cuts of the Fermi-surface segments, which ensures Fermi sea effects are included in the Fermi-liquid picture in order to make the integration tractable Haldane (2004); Wang et al. (2007). Here the Fermi surface of CuMnAs model is not isotropic. We interpret the full OHE as a Fermi sea effect. The transformation of the Fermi sea to Fermi surface calculation does not simplify the full OHE integral. We calculate the orbital Hall conductivity dependence on the Fermi energy $E_{\rm F}$, when $E_{\rm F}$ is in the conduction band. For the tight-binding Hamiltonian of CuMnAs, we obtain the total intrinsic OHE in Fig. 1, showing that the magnitude of both contributions decreases with increasing $E_{\rm F}$ Liu and Culcer (2024); Bhowal and Vignale (2021). Remarkably, the conventional contribution corresponds to only one fraction of the total orbital current, and the quantum correction $\Delta j$ accounts for almost the entire value. We also plot the individual contributions to $\Delta j$ in Fig. 2, showing that in this model the dominant term is $\Delta j_{3}$, which arises from the non-commutativity of the position and velocity operators.
Discussion. We begin with the rationale behind our approach. According to the philosophy of quantum mechanics, the starting point in any calculation is the operator under consideration. This approach has underpinned the entire modern theory of orbital magnetization, which takes the OAM operator as the starting point in the evaluation of the OAM. The equilibrium OAM stems from the inter-band matrix elements of the position and velocity operators, in fact it can be regarded as an inter-band coherence phenomenon in the sense of Ref. Culcer et al. (2017), albeit an equilibrium version of it. Traditionally these inter-band matrix elements were neglected in transport calculations Ashcroft and Mermin (1976); Cardona (1969), and the OAM would have continued to be missed if this neglect persisted. According to this reasoning, the starting point of the orbital current calculation is the orbital current operator, whose matrix elements must be evaluated in their entirety. Such a full quantum mechanical evaluation reveals the presence of several contributions to the orbital current, which affect the composition and magnitude of the orbital Hall effect. For the model system chosen the quantum correction is much larger in magnitude than the conventional value. In CuMnAs the conventional and corrected OHE also differ by a sign, yet there is no guarantee that such a difference exists in general. To determine the sign in principle the OHE can be imaged using the techniques developed by Zeldov, although this remains to be demonstrated Grover et al. (2022).
The equilibrium OAM is expressed purely in terms of off-diagonal matrix elements of the position and velocity operators Sundaram and Niu (1999); Chang and Niu (1995, 1996), hence the unruly position operator is tamed Thonhauser et al. (2005); Ceresoli et al. (2006); Resta (2020). Nevertheless, as our Supplement shows, the presence of only inter-band position and velocity matrix elements in the equilibrium OAM is a result of the equilibrium density matrix being purely diagonal in the band index. In a non-equilibrium scenario the evaluation of the OAM and related quantities such as the orbital current may contain band-diagonal matrix elements of the position and velocity operators.
How is one to interpret the presence of band-diagonal position matrix elements, given that Bloch electrons are delocalized and their position is undefined Yafet (1963); Esteve-Paredes and Palacios (2023)? We recall that the discussion of Bloch electrons’ position is entirely based on an equilibrium picture, and the equilibrium location of an electron in the lattice is indeed undefined. Nevertheless, out of equilibrium an electric field in general displaces an electron’s center of mass away from its equilibrium location, generating a dipole. Focusing on the intrinsic case, in which electric-field induced changes to the density matrix are purely band off-diagonal, this dipole represents a DC version of the inter-band polarization known in non-linear optics Sipe and Shkrebtii (2000); King-Smith and Vanderbilt (1993). This suggests one mechanism behind the OHE as the generation of an inter-band polarization by the applied electric field: the electric field generates a dipole which then rotates about the centre of mass, generating an OAM, which is convected along with the electron. This is the physics behind $\Delta j_{1}$. In incorporating the band-diagonal position matrix elements one must ensure that no part of the final result depends on the equilibrium centre of mass of the electron. Our boundary conditions take this into account.
Building on this insight, if the OAM is regarded as an operator, one must all for this operator to have inter-band matrix elements, and these can contribute to OAM transport, in the same way as the inter-band position and velocity matrix elements are responsible for the equilibrium OAM. Such band off-diagonal matrix elements of the OAM operator reflect inter-band coherence induced by an electric field. It is these matrix elements that are primarily responsible for $\Delta j_{2}$. Next, once the position and velocity are treated as full operators, one must also account for the fact that they do not commute, and this leads directly to $\Delta j_{3}$.
We have focussed on the Bloch picture of electrons in crystals. Different insights into the OAM emerge from the Wannier and Bloch formulations of the modern theory Rhonald Burgos Atencia and Culcer (2024); Cysne et al. (2022). In the Wannier picture, the OAM involves intra-atomic and inter-atomic contributions Pezo et al. (2022, 2023); Busch et al. (2023); Xuan and Quek (2020), yet an exact correspondence between individual terms in the OAM between the two approaches is not always possible Rhonald Burgos Atencia and Culcer (2024). Nevertheless, since an exact correspondence exists between the Bloch and Wannier bases, the fundamental considerations outlined in this work will be applicable to Wannier studies as well.
We stress that our work is concerned with the evaluation of the current, not its definition. In the long run the same fundamental issues will need to be considered for the OHE as for the spin-Hall effect. One is that the orbital Hall current, as defined, can be nonzero in equilibrium in the same way as the spin current Rashba (2003). In fact all the four contributions to $j_{L}$ could be separately nonzero in equilibrium, although they do vanish for all the effective models we have investigated – CuMnAs, bulk topological insulators, and massive Dirac fermions. Secondly, the OAM itself may not be conserved in an electric field, as we recently showed Atencia et al. (2024), and one may eventually need to consider a proper OH current in analogy with the proper spin current. Thirdly, the only observable is the OAM, and the relationship between the current and the OAM accumulation at the boundary will have to be determined. Nevertheless, no matter which approach is taken, the fundamental problem of the evaluation of $j_{L}$ is relevant. In this context, to examine OAM conservation in our chosen example we calculated the torque term $\langle\partial_{t}\hat{L}_{\alpha}\rangle$ for the CuMnAs tight-binding model Eq. (6) following the prescription of Ref. Atencia et al. (2024), and found $\langle\partial_{t}\hat{L}_{\alpha}\rangle=0$. This implies the OAM is conserved in this model even in an electric field.
Figure 3: The intrinsic $j_{L}$, extrinsic $j_{\rm dis}$, and total OHE for CuMnAs, with the same parameters as in Fig. 1.
Disorder. Non-equilibrium effects can be classified as intrinsic and extrinsic depending on whether they originate in the band structure or disorder. Terms to zeroth order in the disorder strength play crucial roles in Hall effects Khaetskii (2006); Gradhand et al. (2010); Ferreira et al. (2014); Cullen and Culcer (2023); Liu and Culcer (2024). In the density matrix formalism such corrections are incorporated into an anomalous driving term Culcer et al. (2017) which results in a correction to $\rho^{mn}_{E{\bm{k}}}$ in Eq. 4. In a recent work we have argued that this disorder correction dominates the OHE in massless Dirac fermion systems Liu and Culcer (2024), where our calculation was based on the conventional evaluation of the OHE. Here we estimate the extrinsic contribution to the full orbital current in the relaxation-time approximation, where the inverse relaxation time serves as a measure of the disorder strength. The extrinsic OHE is zeroth-order in the relaxation time, that is, the same order as the intrinsic OHE Liu and Culcer (2024). The relaxation time approximation does not take into account the electrical field corrected scattering integral Atencia et al. (2022), which is extremely laborious. We note also that the Eq. (6) is a 4-band model, which is considerably more complex than the two-band tilt Dirac model of Ref. Liu and Culcer (2024). Our results are shown in Fig. 3. We find that the extrinsic contribution (i) is much larger than the conventional OHE, in agreement with Ref. Liu and Culcer (2024), and (ii) is of the same order of magnitude as intrinsic contribution to the total $j_{L}$, including the quantum correction.
Conclusion. We have presented a full quantum mechanical evaluation of the orbital current and orbital Hall effect, demonstrating that the existence of a quantum correction that is missed in conventional evaluations. This correction stems from the band-diagonal terms in the position and velocity operators, as well as the non-commutativity of the position and velocity operators. In the tight-binding Hamiltonian of tetragonal CuMnAs the quantum correction is by far the dominant contribution to the OHE and changes its sign.
Acknowledgments. This work is supported by the Australian Research Council Centre of Excellence in Future Low-Energy Electronics Technologies, project number CE170100039. We acknowledge enlightening discussions with Naoto Nagaosa, Guang-Yu Guo, Hsin Lin, Mikhail Titov, and Tatiana Rappoport.